OpenStudy (anonymous):

a straight line passes through the points A(0,3) and B(3,12). Write down its a) gradient; b) equation ii) the line x=3 is the axis of symmetry of triangleABC. Find the coordinates of C.

OpenStudy (aroub):

a)=12-3/3-0 =9/3 =3 b) y=mx+b , m=3 , (0,3) 3=3(0)+b 3=0+b 3=b => y=3x+3